Please solve this using the elimination method.
9x + 2y = 2
21x + 6y = 4

Answer by checkley77(12844)

9x+2y=2 multiply by -3 & add.
21x+6y=4
-27x-6y=-6
---------------
-6x=-2
x=-2/-6
x=1/3 ans.
9*1/3+2y=2
3+2y=2
2y=2-3
2y=-1
y=-1/2 ans.
Proof:
21*1/3+6*-1/2=4
7-3=4
4=4